OpenStudy (anonymous):

f(x) = x-2/2 then what is f^–1(x)

OpenStudy (anonymous):

u hv to find the inverse. do u know how?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oki

OpenStudy (anonymous):

f(x) = -2x - 7 means: What value comes out when you use some value for x. That right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is your f(x) (x-2)/2 ??

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

The inverse means: if we know the value that comes out, can we find the original value of x?

OpenStudy (anonymous):

lets assume that it is

OpenStudy (anonymous):

we artificially use a different variable to make the operation simpler to follow

OpenStudy (anonymous):

So, let y = f(x) = (x-2)/2

OpenStudy (anonymous):

can u find x in terms of y ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oki, lets do it together: y = (x-2)/2 means, 2y = x-2 clear till here?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

2y + 2 = x means, x = 2y + 2

OpenStudy (anonymous):

we rewrite in terms of x: f^(-1)(x) = 2x + 2

OpenStudy (anonymous):

Thats your answer

OpenStudy (anonymous):

we can check too: f^(-1) [f(x)] = f^(-1)(2x+2) = [(2x+2) - 2]/2 = [+2x + 2 - 2]/2 = 2x/2 = x OK for all values of x